# Homogeneous Stellar Atmospheric Parameters and 22 Elemental Abundances for FGK Stars Derived From LAMOST Low-resolution Spectra with DD-Payne

### v1

@@ Abstract
A deep understanding of our Galaxy desires detailed decomposition of its stellar populations via their chemical fingerprints. This requires precise stellar abundances of many elements for a large number of stars. Here we present an updated catalog of stellar labels derived from LAMOST low-resolution spectra in a physics-sensible and rigorous manner with DD-Payne, taking labels from high-resolution spectroscopy as training set. The catalog contains atmospheric parameters for 6.4 million stars released in LAMOST DR9, and abundances for 22 elements, namely, C, N, O, Na, Mg, Al, Si, Ca, Ti, Cr, Mn, Fe, Ni, Sr, Y, Zr, Ba, La, Ce, Nd, Sm, and Eu, for nearly 3.6 million stars with spectral signal-to-noise ratio (SNR) higher than 20. The $[{\mathrm{Fe/H}}]\,$ is valid down to $\simeq-4.0$, while elemental abundance ratios [X/Fe] are mostly valid for stars with $[{\mathrm{Fe/H}}]\,\gtrsim-2.0$. Measurement errors in these labels are sensitive to and almost inversely proportional with SNR. For stars with ${\mathrm{S}NR}>50$, we achieved a typical error of 30 K in $T_{\mathrm{eff}}\,$, 0.07 dex in $\log~{}g$, $\sim 0.05$ dex in abundances for most elements with atomic number smaller than Sr, and 0.1–0.2 dex for heavier elements. Homogenization to the label estimates is carried out via dedicated internal and external calibration. In particular, the non-local thermal equilibrium effect is corrected for the [Fe/H] estimates, the $T_{\mathrm{eff}}\,$ is calibrated to the infrared flux method scale, and the $\log~{}g$ is validated with asteroseismic measurements. The elemental abundances are internally calibrated using wide binaries, eliminating systematic trend with effective temperature. The catalog is publicly available.
@@ 2.1 The LAMOST survey spectra
Fig. 2 shows the sky coverage of the sample stars in Galactic coordinates. The sample stars have a nearly contiguous distribution in the Galactic latitude range $-10^{\circ}<b<60^{\circ}$. Particularly, the survey has a continuous and dense ($\gtrsim 1000$ stars per square degree) sampling for stars in the anti-center disk ($150^{\circ}\lesssim l\lesssim 210^{\circ}$, $|b|<30^{\circ}$) and the M31/M33 vincility fields $(l,b)\simeq(120^{\circ},-20^{\circ})$, as a result of dedicated targeting by the LAMOST Spectroscopic Survey of the Galactic Anti-center (LSS-GAC; Liu et al., 2014; Yuan et al., 2015; Xiang et al., 2017). At higher Galactic latitudes, the number density drops because the intrinsic stellar density of the Galaxy decreases with increasing latitude. The figure also shows a dense spot at the $Kepler$ field $(l,b)\simeq(70^{\circ},15^{\circ})$, as a result of the dedicated LAMOST-$Kepler$ survey (De Cat et al., 2015a; Fu et al., 2020; Zong et al., 2020).
@@ 4.2 Validating $\log~{}g$ with asteroseismology
Accurate and precise $\log~{}g$ derived with Asteroseismology provide a golden calibration source for spectroscopic $\log~{}g$ estimates (e.g., Ren et al., 2016). Light curves from the $Kepler$ mission (Borucki et al., 2010) allow precise determination of asteroseismic frequency for more than nearly 20 thousands red giant stars (e.g. Yu et al., 2018) and a few hundred of main-sequence and subgiant stars (e.g. Mathur et al., 2022). About half of the $Kepler$ stars are observed by LAMOST via a dedicated survey to the $Kepler$ field (De Cat et al., 2015b; Fu et al., 2020).
@@ 4.3 Correcting ${\rm[Fe/H]}$ for NLTE effect
NLTE effects can have a large and complicated impact on the abundance determinations. For neutral iron and similar minority species, these effects are typically larger towards lower metallicities, as well has for hotter stars and for stars of lower surface gravity (e.g., Mashonkina et al., 2011; Bergemann et al., 2012; Lind et al., 2012; Sitnova et al., 2015; Lind & Amarsi, 2024). The abundances for both the APOGEE DR17 and the high-resolution sample of Li et al. (2022) are derived using model spectra computed under LTE assumption. While our determination taking these labels as training set has achieved a good internal precision, the resultant labels may suffer non-negligible systematic errors due to NLTE effects that are not properly considered in the training labels. In this work, we conduct a NLTE correction for the DD-Payne [Fe/H] determination.
@@ 5 The DD-Payne catalog for LAMOST DR9
Finally, we obtain a catalog of 25 stellar labels, including atmospheric parameters $T_{\rm eff}$, $\log\,g$, ${\rm[Fe/H]}$, and $v_{\rm mic}$, as well as elemental abundance ratio [X/Fe], with X being C, N, O, Na, Mg, Al, Si, Ca, Ti, Cr, Mn, Ni, Sr, Y, Zr, Ba, La, Ce, Nd, Sm, and Eu. The atmospheric parameters and abundance ratio for 13 elements, namely C, N, O, Na, Mg, Al, Si, Ca, Ti, Cr, Mn, Ni, and Ba, are valid for 6439,359 stars ( 9245,463 spectra) from LAMOST DR9. While the abundance ratios for Sr, Y, Zr, La, Ce, Nd, Sm, and Eu are provided only for about 3.6 million stars with SNR higher than 20.
For convenience, we provide two separate tables in .FITS format, one for recommended values of the labels, after applying for the calibrations in Sect. 4. The other is for original abundance measurements without calibrations. Table 3 and Table LABEL:table4 present a description of the columns in the these catalogs. The catalog is available in Zenodo, https://zenodo.org/records/15254859, in fits format. The latest version can be found in the NADC, https://nadc.china-vo.org/res/r101600. Future updates to the catalogs, including new application to LAMOST DR12 with the same DD-Payne pipeline set, are presumed to made public via this NADC link.
@@
SNR_u/g/r/i/z median spectral signal-to-noise ratios in u/g/r/i/z bands
@@ 5.3 Stellar elemental abundances
The Figures show a clear s-process enhanced giant star population at the metallicity range of $-1\lesssim[{\mathrm{Fe/H}}]\,\lesssim-0.5$, with ${\rm[X/Fe]}\gtrsim 1$ for many of the s-process elements, such as Y, Zr, Ba, La, Ce, and Nd. Such s-process enhanced stars also occur in the results of dwarfs for Y, Ba, and La. S-process enhanced stars in LAMOST have been studied in previous work utilizing their [Ba/Fe] (Xiang et al., 2019; Norfolk et al., 2019; Zhang et al., 2024). Here our results provide abundances for many more elements, thus allow a more detailed study on their nature. The Zr-enhanced population is especially a strong feature for giant stars, while there is no counterpart of dwarfs in the range of $[{\mathrm{Fe/H}}]\,\gtrsim-1$. This is due to the contribution of a substantial population of S-type stars (Smith & Lambert, 1990; Keenan, 1954) in the giant sample.
For dwarf stars, there is also a Ba-enhanced population with $[{\mathrm{Fe/H}}]\,\gtrsim-0.2$, which has been found to be consequence of stellar internal element transport due to radiative acceleration in intermediate-mass stars (Xiang et al., 2019). For these stars, their [C/Fe], [Mg/Fe], and [Ca/Fe] can be substantially lower, but their [Ni/Fe] can be higher (Michaud et al., 2015). All of these predicted features are seen in our data. Note that as mentioned above (Sect. 3), for many of the elements such as O, Na, Al, Si, Ti, Sr, Y, Ce, Nd, Sm, and Eu, our abundance estimates at the metal-poor end are problematic due to the lack of appropriate training labels. We have marked those regime in the Figures with shaded area. For stars in these shaded regimes, we set a non-zero value in their abundance flags (Sect.  5.1).

### v2

@@ Abstract
A deep understanding of our Galaxy desires detailed decomposition of its stellar populations via their chemical fingerprints. This requires precise stellar abundances of many elements for a large number of stars. Here we present an updated catalog of stellar labels derived from LAMOST low-resolution spectra in a physics-sensible and rigorous manner with DD-Payne, taking labels from high-resolution spectroscopy as training set. The catalog contains atmospheric parameters for 6.4 million stars released in LAMOST DR9, and abundances for 22 elements, namely, C, N, O, Na, Mg, Al, Si, Ca, Ti, Cr, Mn, Fe, Ni, Sr, Y, Zr, Ba, La, Ce, Nd, Sm, and Eu, for nearly 3.6 million stars with spectral signal-to-noise ratio (S/N) higher than 20. The $[{\mathrm{Fe/H}}]\,$ is valid down to $\simeq-4.0$, while elemental abundance ratios [X/Fe] are mostly valid for stars with $[{\mathrm{Fe/H}}]\,\gtrsim-2.0$. Measurement errors in these labels are sensitive to and almost inversely proportional with S/N. For stars with ${\mathrm{S}/N}>50$, we achieved a typical error of 30 K in $T_{\mathrm{eff}}\,$, 0.07 dex in $\log~{}g$, $\sim 0.05$ dex in abundances for most elements with atomic number smaller than Sr, and 0.1–0.2 dex for heavier elements. Homogenization to the label estimates is carried out via dedicated internal and external calibration. In particular, the non-local thermal equilibrium effect is corrected for the [Fe/H] estimates, the $T_{\mathrm{eff}}\,$ is calibrated to the infrared flux method scale, and the $\log~{}g$ is validated with asteroseismic measurements. The elemental abundances are internally calibrated using wide binaries, eliminating systematic trend with effective temperature. The catalog is publicly available.
@@ 2.1 The LAMOST survey spectra
Fig. 2 shows the sky coverage of the sample stars in Galactic coordinates. The sample stars have a nearly contiguous distribution in the Galactic latitude range $-10^{\circ}<b<60^{\circ}$. Particularly, the survey has a continuous and dense ($\gtrsim 1000$ stars per square degree) sampling for stars in the anti-center disk ($150^{\circ}\lesssim l\lesssim 210^{\circ}$, $|b|<30^{\circ}$) and the M31/M33 vicinity fields $(l,b)\simeq(120^{\circ},-20^{\circ})$, as a result of dedicated targeting by the LAMOST Spectroscopic Survey of the Galactic Anti-center (LSS-GAC; Liu et al., 2014; Yuan et al., 2015; Xiang et al., 2017). At higher Galactic latitudes, the number density drops because the intrinsic stellar density of the Galaxy decreases with increasing latitude. The figure also shows a dense spot at the $Kepler$ field $(l,b)\simeq(70^{\circ},15^{\circ})$, as a result of the dedicated LAMOST-$Kepler$ survey (De Cat et al., 2015; Fu et al., 2020; Zong et al., 2020).
@@ 4.2 Validating $\log~{}g$ with asteroseismology
Accurate and precise $\log~{}g$ derived with Asteroseismology provide a golden calibration source for spectroscopic $\log~{}g$ estimates (e.g., Ren et al., 2016). Light curves from the $Kepler$ mission (Borucki et al., 2010) allow precise determination of asteroseismic frequency for more than nearly 20 thousands red giant stars (e.g. Yu et al., 2018) and a few hundred of main-sequence and subgiant stars (e.g. Mathur et al., 2022). About half of the $Kepler$ stars are observed by LAMOST via a dedicated survey to the $Kepler$ field (De Cat et al., 2015; Fu et al., 2020).
@@ 4.3 Correcting ${\rm[Fe/H]}$ for NLTE effect
NLTE effects can have a large and complicated impact on the abundance determinations. For neutral iron and similar minority species, these effects are typically larger towards lower metallicities, as well as for hotter stars and for stars of lower surface gravity (e.g., Mashonkina et al., 2011; Bergemann et al., 2012; Lind et al., 2012; Sitnova et al., 2015; Lind & Amarsi, 2024). The abundances for both the APOGEE DR17 and the high-resolution sample of Li et al. (2022) are derived using model spectra computed under LTE assumption. While our determination taking these labels as training set has achieved a good internal precision, the resultant labels may suffer non-negligible systematic errors due to NLTE effects that are not properly considered in the training labels. In this work, we conduct a NLTE correction for the DD-Payne [Fe/H] determination.
@@ 5 The DD-Payne catalog for LAMOST DR9
Finally, we obtain a catalog of 25 stellar labels, including atmospheric parameters $T_{\rm eff}$, $\log\,g$, ${\rm[Fe/H]}$, and $v_{\rm mic}$, as well as elemental abundance ratio [X/Fe], with X being C, N, O, Na, Mg, Al, Si, Ca, Ti, Cr, Mn, Ni, Sr, Y, Zr, Ba, La, Ce, Nd, Sm, and Eu. The atmospheric parameters and abundance ratio for 13 elements, namely C, N, O, Na, Mg, Al, Si, Ca, Ti, Cr, Mn, Ni, and Ba, are valid for 6439,359 stars ( 9245,463 spectra) from LAMOST DR9. While the abundance ratios for Sr, Y, Zr, La, Ce, Nd, Sm, and Eu are provided only for about 3.6 million stars with S/N higher than 20.
For convenience, we provide two separate tables in .FITS format, one for recommended values of the labels, after applying for the calibrations in Sect. 4. The other is for original abundance measurements without calibrations. Table 3 and Table LABEL:table4 present a description of the columns in the these catalogs. The catalogs are available in Zenodo, https://zenodo.org/records/15254859, in fits format. They can be found in the NADC, https://nadc.china-vo.org/res/r101600. Future updates to the catalogs, including new application to LAMOST DR12 with the same DD-Payne pipeline sets, are presumed to made public via this NADC link.
@@
S/N_u/g/r/i/z median spectral signal-to-noise ratios in u/g/r/i/z bands
@@ 5.3 Stellar elemental abundances
The Figures show a clear s-process enhanced giant star population at the metallicity range of $-1\lesssim[{\mathrm{Fe/H}}]\,\lesssim-0.5$, with ${\rm[X/Fe]}\gtrsim 1$ for many of the s-process elements, such as Y, Zr, Ba, La, Ce, and Nd. Such s-process enhanced stars also occur in the results of dwarfs for Y, Ba, and La. S-process enhanced stars in LAMOST have been studied in previous work utilizing their [Ba/Fe] (Xiang et al., 2020; Norfolk et al., 2019; Zhang et al., 2024). Here our results provide abundances for many more elements, thus allow a more detailed study on their nature. The Zr-enhanced population is especially a strong feature for giant stars, while there is no counterpart of dwarfs in the range of $[{\mathrm{Fe/H}}]\,\gtrsim-1$. This is due to the contribution of a substantial population of S-type stars (Smith & Lambert, 1990; Keenan, 1954) in the giant sample.
For dwarf stars, there is also a Ba-enhanced population with $[{\mathrm{Fe/H}}]\,\gtrsim-0.2$, which has been found to be consequence of stellar internal element transport due to radiative acceleration in intermediate-mass stars (Xiang et al., 2020). For these stars, their [C/Fe], [Mg/Fe], and [Ca/Fe] can be substantially lower, but their [Ni/Fe] can be higher (Michaud et al., 2015). All of these predicted features are seen in our data. Note that as mentioned above (Sect. 3), for many of the elements such as O, Na, Al, Si, Ti, Sr, Y, Ce, Nd, Sm, and Eu, our abundance estimates at the metal-poor end are problematic due to the lack of appropriate training labels. We have marked those regime in the Figures with shaded area. For stars in these shaded regimes, we set a non-zero value in their abundance flags (Sect.  5.1).